OpenStudy (anonymous):

What is the equation of the line passing through (2, -2) and perpendicular to the line y = -x + 3 in slope-intercept form?

OpenStudy (anonymous):

a) y = x b) y = -x c) y = x - 4 d) y = x + 4

OpenStudy (anonymous):

(y-y1)=-1/m(x-x1) formula m=-1 (x1,y1)=(2,-2) substitute

OpenStudy (anonymous):

ans is option d y=x+4

OpenStudy (anonymous):

thank you!!!!